The density of the lightest metal, lithium (Li), is 5.34 3 102 kg/m 3 . Convert the density to g/cm 3


1
Expert's answer
2021-02-04T07:03:50-0500

d(Li) = 5.34×102 kg/m3


Solution:

1 kg = 1000 g

1 m = 100 cm

Hence,

5.34×102 kg/m3 × (1000 g / 1 kg) × (1 m / 100 cm)3 = 5.34×102 × (1000 g / 106 cm3) = 0.534 g/cm3

d(Li) = 0.534 g/cm3


Answer: The density of lithium (Li) is 0.534 g/cm3.